OpenStudy (drummergirl7):

Which equations show the associative property? Select each correct answer. (−1315+(−638))+(−47)=−(1315+638)+(−47)(−1315+(−638))+(−47)=−(1315+638)+(−47) (−289+(−511))+(−1111)=(−1111)+(−289+(−511))(−289+(−511))+(−1111)=(−1111)+(−289+(−511)) −1720+(−18+(−378))=(−1720+(−18))+(−378)−1720+(−18+(−378))=(−1720+(−18))+(−378) (−910+(−13))+(−113)=−910+(−13+(−113))

OpenStudy (drummergirl7):

@mathstudent55

OpenStudy (drummergirl7):

these are fractions

OpenStudy (drummergirl7):

this is my last question

OpenStudy (mathstudent55):

This is similar to the previous problem. See where the numbers remain in the same order, but the parentheses group numbers together differently.

OpenStudy (drummergirl7):

yes

OpenStudy (drummergirl7):

should i link images

OpenStudy (drummergirl7):

plz help

OpenStudy (eiwoh):

Well you should know what the associative property is first. Do you know?

OpenStudy (drummergirl7):

yes i do :)

OpenStudy (eiwoh):

Great know take that and examine each answer. Find out which makes the most sense, and which does not.

OpenStudy (drummergirl7):

can i link screen shots in openstudy?

OpenStudy (drummergirl7):

ok

OpenStudy (mathstudent55):

You just had a problem that is exactly like this one but with different numbers. Go over the previous problem with my explanations until you understand it. Then you should solve this one easily.

OpenStudy (drummergirl7):

ok

OpenStudy (eiwoh):

There is always a "dumb" answer, my teacher always said. Try to see which one that is.

OpenStudy (mathstudent55):

Where do you see the pattern \((a + b) + c = a + (b + c) \) ?

OpenStudy (eiwoh):

Or listen to @mathstudent55. He knows way more than I do lol.

OpenStudy (drummergirl7):

lol. it isnt d

OpenStudy (mathstudent55):

Option A: (−1315 + (−638)) + (−47) = −(1315+638)+(−47) ^ | here you are factoring a -1. this is not it

OpenStudy (eiwoh):

Good, now keep ruling out. For me it's A. What do you think?

OpenStudy (mathstudent55):

Option B: (−289+(−511))+(−1111)=(−1111)+(−289+(−511)) The numbers on both sides are the same and in the same order. Notice on the left side -289 and -511 grouped together. Then -1111 is added. What happens on the right side?

OpenStudy (drummergirl7):

ok

OpenStudy (drummergirl7):

so its a?

OpenStudy (eiwoh):

Totally fanning @mathstudent55

OpenStudy (drummergirl7):

wait its not a and its not d

OpenStudy (drummergirl7):

ikr

OpenStudy (drummergirl7):

so its ether b and c

OpenStudy (drummergirl7):

or both

OpenStudy (eiwoh):

I guess. Is there an option for both?

OpenStudy (drummergirl7):

yea

OpenStudy (mathstudent55):

In option B, the right side has the same numbers grouped differently, so option B is definitely associative property.

OpenStudy (drummergirl7):

what about c :)

OpenStudy (eiwoh):

I gtg but good job @mathstudent55. And @DrummerGirl7 I hope you get the right answer :)

OpenStudy (mathstudent55):

Now see if C and D are similar to A or B.

OpenStudy (drummergirl7):

byeee :)

OpenStudy (drummergirl7):

ok so a is wrong. b is right c is unknown and d is wrong? right?

OpenStudy (drummergirl7):

i hope i get an A

OpenStudy (drummergirl7):

my answer is b and c

OpenStudy (mathstudent55):

Look at C and figure out what is happening with the parentheses between the left side and the right side of the equal sign.

OpenStudy (drummergirl7):

its not correct i had to take a second look

OpenStudy (drummergirl7):

omg computer is about to die!!

OpenStudy (drummergirl7):

what would u say? just b/ or not

OpenStudy (mathstudent55):



OpenStudy (drummergirl7):

ok so it is c!

OpenStudy (drummergirl7):

lets see what i get

OpenStudy (mathstudent55):

So far we know: A: no B: yes C: yes D: we don't know yet

OpenStudy (mathstudent55):

Now look at D. Does D follow the pattern of B and C, or doesn't it?

OpenStudy (drummergirl7):

i... i got a 40% :(

OpenStudy (drummergirl7):

im still giving u a medal tho :)

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

The only change is grouping, so D is also associative.

OpenStudy (drummergirl7):

ok thank u so much

OpenStudy (mathstudent55):

Now we can conclude: A: no B: yes C: yes D: yes

OpenStudy (drummergirl7):

thxx